y=-3x-3z+11
x=-2y-z+17
z=x-y
1
Expert's answer
2012-12-06T09:13:21-0500

y=-3x-3z+11

x=-2y-z+17
z=x-y

y = -3x - 3(x - y) + 11
x = -2y - (x - y) + 17

y = -6x + 3y + 11 -> 2y = 6x - 11
x = -y - x + 17 -> y = -2x + 17 -> 2y = -4x + 34

6x - 11 = -4x + 34
10x = 45

x = 4.5
y = -2*4.5 + 17 = 8
z = 4.5 - 8 = -3.5